# Scaling-aware rating of count forecasts

### v3

@@ Abstract
Forecast quality should be assessed in the context of what is possible in theory and what is reasonable to expect in practice. Often, one can identify an approximate upper bound to a probabilistic forecast’s sharpness, which sets a lower, not necessarily achievable, limit to error metrics. In retail forecasting, a simple, but often unconquerable sharpness limit is given by the Poisson distribution. When evaluating forecasts using traditional metrics such as Mean Absolute Error, it is hard to judge whether a certain achieved value reflects unavoidable Poisson noise or truly indicates an overdispersed prediction model. Moreover, every evaluation metric suffers from precision scaling: Perhaps surprisingly, the metric’s value is mostly defined by the selling rate and by the resulting rate-dependent Poisson noise, and only secondarily by the forecast quality. For any metric, comparing two groups of forecasted products often yields “the slow movers are performing worse than the fast movers” or vice versa, the naïve scaling trap. To distill the intrinsic quality of a forecast, we stratify predictions into buckets of approximately equal predicted value and evaluate metrics separately per bucket. By comparing the achieved value per bucket to benchmarks, we obtain an intuitive visualization of forecast quality, which can be summarized into a single rating that makes forecast quality comparable among different products or even industries. The thereby developed scaling-aware forecast rating is applied to forecasting models used on the M5 competition dataset as well as to real-life forecasts provided by Blue Yonder’s Demand Edge for Retail solution for grocery products in Sainsbury’s supermarkets in the United Kingdom. The results permit a clear interpretation and high-level understanding of model quality by non-experts.
@@
1 Introduction
@@
Forecasting is an application of Information Technology. Just like any other technology – think of the speed of rockets or the efficiency of power plants – forecasting performance is constrained by fundamental natural bounds. This aspect is often downplayed in practice, and the unachievability of a perfect deterministic forecast is taken as witness of forecasting problems (Bower, 2023). However, even assuming perfect input data quality, what a forecast can achieve in terms of precision and accuracy is not defined by its subscriber’s wishes and its creator’s skills but mainly by noise: “Forecasts characterize and reduce but generally do not eliminate uncertainty”, as succinctly distilled by Gneiting et al. (2007). For a meaningful forecasting evaluation, it is indispensable to quantify the level of uncertainty that is ideally reachable under a given set of available covariates and put the achieved forecasting performance into this context. In the words of Gneiting and Katzfuss (2014), “probabilistic forecasting aims to maximize the sharpness of the predictive distributions, subject to calibration, on the basis of the available information set” . That is, a probabilistic forecast should make the strongest, yet true statement possible.
Metrics that are used in business and operations, such as Mean Absolute Error (MAE) and their normalized variants (errors divided by the mean observation) provide a quantitative evaluation of the forecast, but do not answer to which extent the forecast matches the statistical ideal of maximal sharpness under calibration. We argue below that even skilled and experienced experts cannot develop a reliable intuition that allows them to satisfactorily resolve that judgmental question (“is the forecast good?”) based on traditional metrics. We therefore propose an alternative processing, visualization and rating that allows setting achieved values of traditional metrics into a context to reflect the quality of the forecast.
Forecasting professionals have to handle business stakeholders who sometimes set overly ambitious goals that are driven by gut feeling instead of quantitatively corroborated benchmarks. The reason for the confidence with which forecasting goals are set is that metrics such as percentage errors seem to speak for themselves: 5% appears to be good, 20% fair, and 70% unacceptable. As we show below, such assessment is meaningless, for every metric, unless the ideally achievable distribution of targets (for example, Poissonian) and the scale are set: Normalized and absolute errors, even for ideal forecasts, depend strongly on the predicted value itself, since a forecast’s variance almost always scales with the predicted mean.
@@ 1.1 Precision scaling
There has been much debate on the usage of the “correct” evaluation metric (Davydenko and Fildes, 2014; Hyndman and Koehler, 2006; Gneiting, 2011; Fildes et al., 2022; Wheatcroft, 2022; Petropoulos, F. et al., 2022; Kolassa, 2016; Hewamalage et al., 2023), and on the relevance of this very question (Koutsandreas et al., 2021; Robette, 2023; Kolassa, 2023b). In this paper, we tackle a problem that all metrics share when they are applied to count forecasts: They achieve finite values even in the ideally sharp case and exhibit non-trivial precision scaling. By precision scaling, we refer to the observation that any metric evaluated on a group of forecasted mean values of about 1 (which retailers call “slow-movers”) averages to a different value than for a group of forecasted mean values of about 100 (the “fast-movers”). This difference is often not due to the forecast being systematically “better” or “worse” in the slow- or fast-moving regimes in the sense that certain systematic omissions of features or certain mistakes were done in one or the other regime, but mainly because of the heteroscedasticity of counting distributions: The variance of distributions that describe non-negative integer counts depends on their expectation value. For example, the variance of the Poisson distribution is equal to its mean. This is in stark contrast to continuous quantities, for example, temperature, for which one can envisage a truly homoscedastic forecast, such as a normal distribution with constant, temperature-independent variance.
@@
For the Poisson distribution, absolute errors increase and normalized errors (absolute errors divided by the mean observation) decrease with increasing rate. This is why current approaches such as the one described by Bower (2023), a “basic heat map or listing of worst-case forecast error – at either the item or product-family level” are heavily infected by the naïve scaling trap, the mis-interpretation of a symptom caused by precision scaling as a genuine signal of model quality. We will see below upon both an academic example (a baseline model on the the M5 dataset) and upon a real-world example (the forecasts provided by Blue Yonder using their Demand Edge for Retail solution to UK retailer Sainsbury’s) that the naïve scaling trap is the prevalent effect that one observes when evaluating different groups against each other. We thus need to make the forecast judgment procedure aware of the underlying ideal distribution by judging each bucket of similar predictions separately. For a Poisson-limited forecast, a normalized error (mean absolute error divided by mean observation) of about 70% for a slow-mover of selling rate of about 1 per day can be judged as “excellent”, while a 20% normalized error for a fast-mover that sells about 100 pieces per day is disappointing. Similar velocity-dependent behavior can be expected in other contexts.
@@ 1.2 Forecast comparison use cases
Model comparison is the least problematic application of evaluation metrics: Precision scaling artifacts that affect a metric affect the two (or more) competing models in similar ways. When a business-relevant overall metric, which quantifies the business impact (Robette, 2023), has improved on a given fixed dataset for a new model, aside from pathological edge cases, one can typically conclude that the newer model yields better performance. Nevertheless, it is helpful to be given additional information such as at which selling rates the improvement has occurred, and whether that improvement was evenly applied across all velocities, or favors the slow- or fast-movers.
@@ 1.3 Judgement of forecasts
Our method can be applied to any forecasting metric. In a nutshell, it consists of segregating the set of predictions and actuals into buckets that share approximately the same prediction value. All prediction/actual pairs in one bucket then naturally come with the approximately same expectation value for any forecasting metric. For each bucket, we establish which is the theoretical lower bound to the forecasting metric, and which reference values of that metric can still be considered “excellent”, “good”, “OK”, “fair”, “insufficient” and “unacceptable”. Reference values are scaled to allow transfer between different buckets so the grades only need to be defined for one reference bucket, which we will elaborate in Section 3.3. The forecasting metric is then evaluated for each bucket on the actual observations and compared to the reference values to provide a rating of the quality in the respective bucket. The thereby achieved ratings per bucket are re-aggregated to provide an overall rating (“good”, “OK”, “fair”…), or reference rating values for the overall achieved metric (“the overall MAE amounts to 5.4, which is ‘good‘”).
@@ 2.1 A simple approximate upper sharpness bound: The Poisson distribution
Establishing the sharpest possible forecast under given circumstances can be challenging, but one can nevertheless sometimes find approximate benchmarks. In our focus area of non-personalized grocery retail, it is helpful to employ the Poisson distribution as a simple and intuitive working horse (Tichy, 2023c). Let us recapitulate the prerequisites that go into this assumption: Consider the number of sold items $s$ of a product in a store on a given day, for a well attended supermarket with many ($N>100$) customers per day. Our forecast predicts the total number of customers $N$, and the average probability that they buy a given product, which we assume to be small, $p<0.05$. Moreover, no customer buys two or more pieces of the same item. We then deal with $N$ independent Bernoulli-events with success probability $p$, such that the probability to sell $s$ items on that day is binomial,
@@ 2.2 Corrections to the Poisson distribution
In practice, the Poisson assumption layed out above is never perfectly fulfilled. Several effects both increase and decrease the width of the distribution that one could ideally achieve to predict.
@@
The assumption that only one item is bought per customer – underlying the Bernoulli process – is also slightly broken in practice: At Sainsbury’s UK, about 87.5% of sold items are bought alone, 9.8% are bought in pairs of two pieces, about 1.5% in triplets of three pieces, four or more pieces are also ocurring. The variance of the resulting distribution is thereby increased. In certain areas of retail, a buy of many pieces of the same item is not the exception but the rule: In hardware stores, bathroom tiles of a certain kind are not sold at all on most days, until someone wants to refurbish their bathroom and buys a large number. For such products, the Poisson assumption is clearly unsuitable. In grocery retail, however, the effect amounts to a manageable correction.
@@
A priori, it is difficult to estimate the strength of these effects. For the datasets that we employed – M5 and Sainsbury’s – we will see, ex post, that the Poisson limit remains a meaningful first-order benchmark, and that the effects mentioned in this section can be considered “second-order terms”, well behind precising scaling. Quantifying their impact more precisely is certainly desirable, and we plan to do so in future works. In general, if a more suitable benchmark distribution is available, which may include effects such as overdispersion, zero-inflation, and right censoring, our method can easily be adapted, mutatis mutandis.
Hence, to focus on our main contribution, the scaling-aware rating of forecasts across selling rates, our working assumption is that a Poisson forecast constitutes a “best possible” prediction.
@@ 2.4.1 Loss functions on point estimates
where the optimal point estimator for MAE, the median of the predicted distribution, replaces the mean $r_{j}$ (Schwertman et al., 1990; Gneiting, 2011). That is, the forecaster is aware that their forecast $r_{j}$ is not a universal best guess that can be fed into any metric, but that, depending on the chosen metric, an appropriate point estimate needs to be extracted (Kolassa, 2020). MAE naturally takes larger values for larger predictions, a first attempt towards a scale-free metric is the dimensionless normalized MAE, often called Weighted Mean Absolute Percentage Error (WMAPE),
@@
The still-popular Mean Absolute Percentage Error (MAPE) comes with many flaws (Hyndman and Koehler, 2006; Kolassa and Martin, 2011; Kolassa and Schütz, 2007). In particular, it is undefined for observations $s=0$, whose treatment heavily affects the achieved MAPE-values (Kolassa, 2023a). These problems are complemented by the particularly complicated precision scaling of MAPE under the Poisson distribution described in Tichy (2023a), in which a simple removal of events with $s=0$ is assumed.

### v4

@@ Abstract
Forecast quality should be assessed in the context of what is possible in theory and what is reasonable to expect in practice. Often, one can identify an approximate upper bound to a probabilistic forecast’s sharpness, which sets a lower, not necessarily achievable, limit to error metrics. In retail forecasting, a simple, but often unconquerable sharpness limit is given by the Poisson distribution. When evaluating forecasts using traditional metrics such as Mean Absolute Error, it is hard to judge whether a certain achieved value reflects unavoidable Poisson noise or truly indicates an over-dispersed prediction model. Moreover, every evaluation metric suffers from precision scaling: Perhaps surprisingly, the metric’s value is mostly defined by the selling rate and by the resulting rate-dependent Poisson noise, and only secondarily by the forecast quality. For any metric, comparing two groups of forecasted products often yields “the slow movers are performing worse than the fast movers” or vice versa, the naïve scaling trap. To distill the intrinsic quality of a forecast, we stratify predictions into buckets of approximately equal predicted value and evaluate metrics separately per bucket. By comparing the achieved value per bucket to benchmarks, we obtain an intuitive visualization of forecast quality, which can be summarized into a single rating that makes forecast quality comparable among different products or even industries. The thereby developed scaling-aware forecast rating is applied to forecasting models used on the M5 competition dataset as well as to real-life forecasts provided by Blue Yonder’s Demand Edge for Retail solution for grocery products in Sainsbury’s supermarkets in the United Kingdom. The results permit a clear interpretation and high-level understanding of model quality by non-experts.
@@
1 Introduction: The unsolved problem of forecast rating under precision scaling
@@
Forecasting is an application of Information Technology. Just like any other technology – think of the speed of rockets or the efficiency of power plants – forecasting performance is constrained by fundamental natural bounds. This aspect is often downplayed in practice, and the unattainability of a perfect deterministic forecast is taken as witness of forecasting problems (Bower, 2023). However, even assuming perfect input data quality, what a forecast can achieve in terms of precision and accuracy is not defined by its subscriber’s wishes and its creator’s skills but mainly by noise: “Forecasts characterize and reduce but generally do not eliminate uncertainty”, as succinctly distilled by Gneiting et al. (2007). A probabilistic forecast should make the strongest, yet true statement possible. In the words of Gneiting and Katzfuss (2014), “probabilistic forecasting aims to maximize the sharpness of the predictive distributions, subject to calibration, on the basis of the available information set”.
1.1 Problem statement
For a meaningful forecasting evaluation, it is therefore indispensable to quantify the uncertainty that is ideally reachable under a given set of available covariates and to view the achieved forecasting performance in that context. This is, however, not at all what is done in practice. Metrics that are used in business and operations, such as Mean Absolute Error (MAE) and their normalized variants (errors divided by the mean observation, discussed in detail in Section 2.4) do not answer at all to which extent the forecast matches the statistical ideal of maximal sharpness under calibration. Forecasters routinely ask and answer “what MAE do I achieve?”, but, in our experience, they don’t ask “what MAE could I ideally achieve?”. The answer to the first question, however, cannot be interpreted properly without answering the second!
As we will show below, even skilled experts cannot develop a reliable intuition that allows them to intuitively answer the second question, due to precision scaling: The possibly achievable MAE depends strongly on the forecasted value itself. When precision scaling is ignored, one falls into what we name the “naïve scaling trap” – an ostensible effect in forecast quality is not real, but only a consequence of precision scaling.
As a consequence, the problem of rating forecast quality beyond relative comparisons (one forecast is better than some competing method, when evaluated on a fixed set of data) remains largely open. Comparisons across subsets of data (e.g. comparisons across product groups), which are being routinely performed by directly comparing scaling-infected raw metrics (Bower, 2023), are doomed to be infected by the naïve scaling trap and lead to erroneous conclusions (“the slow-moving shoes are performing worse than the fast-moving dairy products”).
Here, we solve the problem of forecast rating by a well-defined sequence of data processing and visualization techniques. These account for scaling and put the achieved values of traditional metrics into the context of what is ideally possible. By explicitly answering the second question (“what MAE could I ideally achieve”), the user then obtains a vivid picture of the true quality of a forecast. This allows them to focus improvement efforts on those subsets of data that truly bear improvement potential instead of those that only ostensibly perform weakly.
In a nutshell, our method consists of segregating the set of predictions and actuals into buckets that share approximately the same prediction. All predictions in one bucket then naturally come with the approximately same expectation value for any forecasting metric. For each bucket, we establish which is the theoretical lower bound to the forecasting metric, and which reference values of that metric can still be considered “excellent”, “good”, “OK”, “fair”, “insufficient” and “unacceptable”. Reference values are scaled to allow transfer between different buckets so the grades only need to be defined for one reference bucket, which we will elaborate in Section 3.3. The forecasting metric is then evaluated for each bucket on the actual observations and compared to the reference values to provide a rating of the quality in the respective bucket. The thereby achieved ratings per bucket are re-aggregated to provide an overall rating (“good”, “OK”, “fair”…), or reference rating values for the overall achieved metric (“the overall MAE amounts to 5.4, which is ‘good‘”). Thereby, we answer the question “what MAE could I ideally achieve?” and set the measured value into that context.
@@ 1.2 Precision scaling
Forecasting professionals have to handle business stakeholders who sometimes set overly ambitious goals that are driven by gut feeling instead of quantitatively corroborated benchmarks. The reason for the confidence with which forecasting goals are set is that metrics such as percentage errors seem to speak for themselves: 5% appears to be good, 20% fair, and 70% unacceptable. As we show below, such assessment is meaningless, for every metric, unless the ideally achievable distribution of targets (for example, Poissonian) and the scale are set: Normalized and absolute errors, even for ideal forecasts, depend strongly on the predicted value itself, since the variance of distributions that describe non-negative integer counts depends on their expectation value. For example, the variance of the Poisson distribution is equal to its mean. This is in stark contrast to continuous quantities, for example, temperature, for which one can envisage a truly homoscedastic forecast, such as a normal distribution with constant, temperature-independent variance.
This common but often overlooked property is precision scaling. As a consequence, any metric evaluated on a group of forecasted mean values of about 1 (which retailers call “slow-movers”) averages to a different value than for a group of forecasted mean values of about 100 (the “fast-movers”). This difference is often not due to the forecast being systematically “better” or “worse” in the slow- or fast-moving regimes in the sense that certain systematic omissions of features or certain mistakes were done in one or the other regime, but mainly because of the heteroscedasticity of counting distributions.
@@
1.3 Literature background
There has been much debate on the usage of the “correct” evaluation metric (Davydenko and Fildes, 2014; Hyndman and Koehler, 2006; Gneiting, 2011; Fildes et al., 2022; Wheatcroft, 2022; Petropoulos, F. et al., 2022; Kolassa, 2016; Hewamalage et al., 2023). From a practitioner’s point of view, it is important to bear in mind that forecast quality is not a goal by itself, but only the means to achieve the optimal business decision, which makes forecast quality unimportant in certain cases (Koutsandreas et al., 2021; Robette, 2023; Kolassa, 2023b).
In general, there is a gap between the statistics state of the art and what is done in practice. Statisticians naturally see forecasts as probabilistic, falsifiable statements, and use the tools of probability theory to evaluate them (Gneiting and Katzfuss, 2014). This approach leads, among others, to the notions of calibration and sharpness (Gneiting and Raftery, 2004; Gneiting et al., 2007) and optimal point forecasts (Gneiting, 2011). Many superficially surprising or paradoxical properties of metrics can then be understood best when they are investigated in terms of those concepts (Gneiting, 2011; Kolassa, 2020).
The effect that we focus on here, precision scaling, is due to the scaling properties of the sharpest possible distribution in a given forecasting context. To our knowledge, precision scaling has not been investigated so far, and computing expectation values of metrics (or other statistics) under a certain distribution is not common in forecasting practice, in great contrast to physics, for example.
Many dimensionless metrics have been proposed. For example, when the popular Mean Absolute Error is divided by the mean observation (Kolassa and Schütz, 2007), one obtains the dimensionless normalized MAE. Other attempts to obtain scale-free measures are MASE (mean absolute scaled error) (Hyndman and Koehler, 2006). Mean Absolute Percentage Error (MAPE) remains popular due to its superficial simplicity, despite strong criticism (Kolassa and Martin, 2011; Tichy, 2023a; Kolassa, 2023a). Among other problems, MAPE is unbounded when over-forecasting, but bound by 100% when under-forecasting. This asymmetry lets it favor biased point forecasts. The symmetric MAPE introduced by Makridakis (1993) resolves this asymmetry only ostensibly, as pointed out by Goodwin and Lawton (1999). Other adaptations of MAPE that aim at overcoming its asymmetry use the logarithm of the forecast-to-actual ratio (Tofallis, 2013). All these dimensionless metrics remain scale-dependent in the sense that they assume different values for slow and for fast velocities under a maximally sharp and calibrated forecast.
The prevalence of precision scaling in all distribution-agnostic metrics is not at all surprising: Precision scaling is a property of how the family of sharpest distributions behaves when the parameter reflecting the expected value is changed, not a property of the metric. Any attempt to judge the achieved values of metrics therefore needs to explicitly incorporate the sharpest possible distribution.
Ex post, the lack of interest in precision scaling is surprising, given the enormous impact that this effect has on forecasting KPIs and the simplicity of computing expectation values. It exemplifies the wide gap between statisticians and practitioners, which we hope to help close.
1.4 Precision scaling in the Poisson distribution
@@
For the Poisson distribution, absolute errors increase and normalized errors (absolute errors divided by the mean observation) decrease with increasing Poisson process rate. This is why current approaches such as the one described by Bower (2023), a “basic heat map or listing of worst-case forecast error – at either the item or product-family level” are heavily infected by the naïve scaling trap, the mis-interpretation of a symptom caused by precision scaling as a genuine signal of model quality. We will see below upon both an academic example (a baseline model on the the M5 dataset, Section 4) and upon a real-world example (the forecasts provided by Blue Yonder using their Demand Edge for Retail solution to UK retailer Sainsbury’s, Section 5) that the naïve scaling trap is the prevalent effect when evaluating different groups against each other. We thus need to make the forecast judgment procedure aware of the underlying ideal distribution by judging each bucket of similar predictions separately. For a Poisson-limited forecast, a normalized error (mean absolute error divided by mean observation) of about 70% for a slow-mover of selling rate of about 1 per day can be judged as “excellent”, while a 20% normalized error for a fast-mover that sells about 100 pieces per day is disappointing. Similar velocity-dependent behavior also prevails in other contexts.
@@ 1.5 Forecast comparison use cases
Model comparison is the least problematic application of evaluation metrics: Precision scaling artefacts that affect a metric affect the two (or more) competing models in similar ways. When a business-relevant overall metric, which quantifies the business impact (Robette, 2023), has improved on a given fixed dataset for a new model, aside from pathological edge cases, one can typically conclude that the newer model yields better performance. Nevertheless, it is helpful to be given additional information such as at which selling rates the improvement has occurred, and whether that improvement was evenly applied across all velocities, or favors the slow- or fast-movers.
@@
1.6 Outline
@@ 2.1 A simple approximate upper sharpness bound: The Poisson distribution
Establishing the sharpest possible forecast under given circumstances can be challenging, but one can nevertheless sometimes find approximate benchmarks. In non-personalized grocery retail, it is helpful to employ the Poisson distribution as a simple and intuitive working horse (Tichy, 2023c). Let us recapitulate the prerequisites that go into the Poisson assumption: Consider the number of sold items $s$ of a product in a store on a given day, for a well attended supermarket with many ($N>100$) customers per day. Our forecast predicts the total number of customers $N$, and the average probability that they buy a given product, which we assume to be small, $p<0.05$. Moreover, no customer buys two or more pieces of the same item. We then deal with $N$ independent Bernoulli-events with success probability $p$, such that the probability to sell $s$ items on that day is binomial,
@@ 2.2 Corrections to the Poisson distribution
In practice, the Poisson assumption laid out above is never perfectly fulfilled. Several effects both increase and decrease the width of the distribution that one could ideally achieve to predict.
@@
The assumption that only one item is bought per customer – underlying the Bernoulli process – is also slightly broken in practice: At Sainsbury’s UK, about 87.5% of sold items are bought alone, 9.8% are bought in pairs of two pieces, about 1.5% in triplets of three pieces, four or more pieces are also occurring. The variance of the resulting distribution is thereby increased. In certain areas of retail, a buy of many pieces of the same item is not the exception but the rule: In hardware stores, bathroom tiles of a certain kind are not sold at all on most days, until someone wants to refurbish their bathroom and buys a large number. For such products, the Poisson assumption is clearly unsuitable. In grocery retail, however, the effect amounts to a manageable correction.
@@
A priori, it is difficult to estimate the strength of these effects. For the datasets that we employed – M5 and Sainsbury’s – we will see, ex post, that the Poisson limit remains a meaningful first-order benchmark, and that the effects mentioned in this section can be considered “second-order terms”, well behind precision scaling. Quantifying their impact more precisely is certainly desirable, and we plan to do so in future works. In general, if a more suitable benchmark distribution is available, which may include effects such as over-dispersion, zero-inflation, and right censoring, our method can easily be adapted, mutatis mutandis.
Hence, to focus on our main contribution, the scaling-aware rating of forecasts across selling rates, our working assumption is that a Poisson forecast constitutes the best possible, that is, the sharpest possible, prediction.
@@ 2.4.1 Loss functions on point estimates
where the optimal point estimator for MAE, the median of the predicted distribution, replaces the mean $r_{j}$ (Schwertman et al., 1990; Gneiting, 2011). That is, the forecaster is aware that their forecast $r_{j}$ is not a universal best guess that can be fed into any metric, but that, depending on the chosen metric, an appropriate point estimate needs to be extracted (Kolassa, 2020). MAE naturally takes larger values for larger predictions, a first attempt towards a scale-free metric is the dimensionless normalized MAE (Kolassa and Schütz, 2007), often called Weighted Mean Absolute Percentage Error (WMAPE),
@@
The still-popular Mean Absolute Percentage Error (MAPE) comes with many flaws (Hyndman and Koehler, 2006; Kolassa and Martin, 2011; Kolassa and Schütz, 2007). In particular, it is undefined for observations $s=0$, whose treatment heavily affects the achieved MAPE-values (Kolassa, 2023a). These problems are complemented by the particularly complicated precision scaling of MAPE under the Poisson distribution described by Tichy (2023a), in which a simple removal of events with $s=0$ is assumed.